What is the largest number that can be represented using eight bits? Select one: a. 255 O b. 256 c. 128 O d. 127
Answer #1

Solution:

Largest 8 bit number is eight 1s i.e. 11111111.

Now 11111111 is 2^0 * 1 + 2^1 * 1 + 2^2 * 1 + 2^3 * 1 + 2^4 * 1 + 2^5 * 1 + 2^6 * 1 + 2^7 * 1 = 1 + 2 + 4 + 8 + 16 + 32 + 64 + 128 = 255.

Note: To convert binary to decimal we multiply the bit value with 2 raised to the power (where power is 0 for bit at extreme right and keeps incrementing by 1 moving towards left).

Therefore, the answer is 255.

The correct option is Option a (255).